OpenStudy (anonymous):

Help! Choose the point-slope form of the equation below that represents the line that passes through the point (6, -3) and has a slope of 1/2. y-6=1/2x+3) y=1/2x-6 y+3=1/2(x-6) x-2y=12

OpenStudy (kira_yamato):

\[y - y_0 = m(x - x_0)\] Just rearrange the terms to get \[y = mx - mx_0 + y_0\] Slope = m \[c = y_0 - mx_0\]

OpenStudy (anonymous):

I don't understand. ._.

OpenStudy (kira_yamato):

m is the slope, and the given point has coordinates \[(x_0 , y_0)\]

OpenStudy (anonymous):

Oh, okay. Thank you.